A solenoid of length L = 10 m carrying a current I = 9 A has a magnetic field of 2 x 10-5 T. Calculate the number of turns. 
Answer #1

B=Magnetic field due to solenoid= unI

u=Permeability of space

n=N/L=Number of Turns per unit length

N=Number of turns

L=Length

I=Current flowing

B= uNI/L

2*10-5= 1.256*10-6*N*9/10

N=17.69 turns